OpenStudy (anonymous):

A scale on a map shows that 2 inches = 25 miles. Part A: How many inches on the map represents 60 miles? Part B: How many miles is represented by 1/4 inch on the map?

OpenStudy (jdoe0001):

\(\large { \begin{array}{ccllll} inches&miles \\\hline\\ 2&25\\ x&60 \end{array}\implies \cfrac{2}{x}=\cfrac{25}{60}\impliedby \textit{solve for "x"} \\ \quad \\ \begin{array}{ccllll} inches&miles \\\hline\\ 2&25\\ \frac{1}{4}&x \end{array}\implies \cfrac{2}{\frac{1}{4}}=\cfrac{25}{x}\impliedby \textit{solve for "x"} }\)

OpenStudy (anonymous):

Um what

OpenStudy (anonymous):

Do you still need help?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i will medal

OpenStudy (anonymous):

okay so we know that 2 inches = 25 miles right

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

You're pretty much setting up a proportion. 2 inches / 25 miles = x / 60. This also means that 25 x = 120 through cross multiplication. \[\frac{ 2 }{ 25 } = \frac{x}{60}\] That's the fraction that jdoe0001 created for you.

OpenStudy (anonymous):

i still don't get it i was never good at cross multiplication

OpenStudy (anonymous):

The cross multiplication turns into 25x = 120. Solving for x will answer the first part of your question. For your second part, you say 2x = 6.25 (since 25 x 1/4 = 6.25) and solve for x there.

OpenStudy (anonymous):

so your saying that 25 times x equals 120

OpenStudy (anonymous):

Yes. I am also saying that 2 times x equals 6.25 for your second question.

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Sure thing.